1. Question Details Q4 P1 [4006861] - A climber whose mass is 53 kg hangs from a rope 11 m below a ledge. What is the magnitude of the tension in the rope? tension Later, his friend pulls him up with a constantly increasing speed that is initially 0 m/s when he is at the bottom and 5 m/s when he reaches the ledge. It takes 70 s to pull him up. What is the y-component of the rate of change of momentum of the dt Now what is the magnitude of the tension in the rope while he is being pulled up? tension-
Answer #1

here,

mass , m = 53 kg

length , l= 11 m

a)

the tension in the rope , T = m * g

T = 519.9 N

b)

time taken , t = 70 s

final speed , v = 5 m/s

let the accelration be a

v = u + a * t

5 = 0 + a * 70

solving for a

a = 0.071 m/s^2

the rate of change of momentum = m * a = 53 * 0.071 = 3.79 kg.m/s^2

c)

let the magnitude of tension in the rope be T'

T' - m * g = m * a

T' = 53 * ( 9.81 + 0.071) N

T' = 523.6 N